A pilot sets an air speed and heading of 575. km/h
Answer #1

Speed of plane = -575cos(36.6)i - 575sin(36.6)j = -461.62i - 342.83j

a) Velocity of plane wrt air = (-461.62i - 342.83j) - (58i + 58j) = -519.62i -400.83j

b) Velocity of plane wrt station = (-461.62i - 342.83j) + (58i + 58j) = -403.61i - 284.83j